OpenStudy (anonymous):

Coffee cost $12 per case and Tea cost $8 per case. If an Order comes in for a total of 250 cases for $2,600, what was the specific number of cases of Tea?

OpenStudy (anonymous):

x: number of cases of coffee y: number of cases of tea x + y =250 12 x + 8 y = 2600 x + y =250 ( multiply by -12) -12x - 12y = -3000 12x + 8 y = 2600 -4y=-400 y= 100